# LLM-enhanced Air Quality Monitoring Interface via Model Context Protocol

### v1

@@ I Introduction
With the growing emphasis on healthy living, air quality monitoring has become a significant area of research. Numerous studies have shown that indoor air pollutants can significantly affect human health [pillarisetti2022indoor]. Particulate Matter (PM) suspended in the air, particularly PM2.5, has been proven to be directly associated with various health risks including cardiovascular diseases, respiratory infections, and even cancer [world2021global]. Therefore, developing air quality monitoring systems that are real-time, accurate, and easy to understand is crucial for enhancing public health awareness and formulating effective environmental policies.
Existing air quality monitoring systems still face numerous challenges in practical applications. Traditional monitoring stations, while capable of providing high-precision data, have high deployment costs and limited coverage [colleaux2025air], resulting in insufficient spatial and temporal resolution of data [cromar2019air]. More importantly, these systems primarily present data through complex charts and indicators, making it difficult for ordinary users without professional backgrounds to intuitively understand the meaning behind the data and make informed decisions accordingly. Furthermore, traditional systems generally have poor interactivity, preventing users from querying specific questions of interest in a natural and direct manner, thereby limiting the practical value of data and public engagement [saini2020comprehensive].
@@
Despite the tremendous potential of LLMs in the field of data analysis, reliably integrating them with real-time environmental monitoring data remains an under-explored research topic. Traditional LLM applications often suffer from “hallucinations” due to lack of real-time contextual knowledge, providing inaccurate or outdated information [huang2025survey]. Unlike traditional retrieval-based methods, our system leverages the Model Context Protocol (MCP) to allow LLMs to securely access real-time IoT data, ensuring grounded, context-aware responses [hou2025model]. Figure 1 illustrates the overall architecture of our proposed system. Our main contributions can be summarized as follows:
@@ II Related Work
Air Quality Monitoring and Accessibility: Traditional air quality monitoring systems have matured in terms of sensor networks and data acquisition. Saini et al. provided a comprehensive overview of indoor air quality monitoring systems, documenting advances in sensor network design, pollutant detection, and data collection mechanisms [saini2020comprehensive]. These systems often achieve impressive technical performance. However, Gupta et al. shown that the information presented by many digital platforms is not always interpretable to non-expert audiences, revealing a significant gap between data acquisition and user-friendly communication of results [gupta2022human]. Morawska et al. emphasized the rapid proliferation of low-cost sensors and community-driven monitoring networks, which greatly improve spatial and temporal resolution, yet, the sheer volume and heterogeneity of data can overwhelm typical interfaces, reducing their effectiveness for general users [morawska2018applications]. Recently, Aguado et al. evaluated low-cost indoor air quality sensors in real-world settings for CO2 and particulate matter, and while the devices supported trend tracking, the study underscores the need for intuitive feedback systems that help non-experts understand what the readings mean for health and behavior [aguado2025verification]. Taken together, this body of work shows that although sensor technology has matured and become more affordable, the critical challenge remains: designing interfaces that can translate dense, technical measurements into actionable insights that non-technical users can easily grasp and act upon.
@@
LLMs for Conversational Interfaces: Recent advances in LLMs have introduced new opportunities to make complex data more accessible through conversational interfaces. By allowing natural language queries, LLMs lower the barrier for non-technical users. However, In two separate studies, Dam et al. and Huang et al. both emphasized a central challenge: the propensity of LLMs for hallucination, where models produce fluent but factually incorrect responses when operating without sufficient context [dam2024complete, huang2025survey]. To mitigate this, Brown et al. demonstrated that directly injecting sensor data into the model’s prompt could enable context-aware summaries and answers about specific datasets [brown2020language]. While effective at small scale, their approach is constrained by the finite context windows of current LLMs. More recently, Islam et al. proposed a prompt ensemble strategy that improved reliability of LLMs response by aggregating outputs from multiple prompts, offering a promising pathway to reduce hallucination and enhance robustness in domain-specific applications [islam2025llm].
@@
Retrieval Augmented Generation for Scalability: The scalability limitations of direct context injection have been well established, most notably through the “lost-in-the-middle” phenomenon, where model performance degrades when long sequences are provided [liu2023lost]. To address this issue, Gao et al. introduced Retrieval-Augmented Generation (RAG) as a dominant framework [gao2023retrieval]. RAG operates by retrieving a targeted subset of relevant information from a larger knowledge base and passing only this filtered context to the LLM. While this approach has proven effective across a wide range of applications, Barnett et al. pointed out that its success is highly sensitive to the quality of the retrieval process, including choices around chunking strategy and parameter tuning [barnett2024seven]. Furthermore, RAG is not ideally suited for structured, time-series data such as environmental sensor streams, where semantic chunking fails to capture the inherent temporal dependencies in the data.
@@
Toward Agentic Systems with MCP: Building on the limitations of retrieval-based methods, our work adopts an active, agentic framework. Rather than having the system predefine which information to supply, the decision is delegated directly to the model. To enable this, Anthropic et al. introduced the Model Context Protocol (MCP), which allows LLMs to autonomously call a predefined set of secure tools to interact with backend systems [anthropic2024mcp]. This approach provides two key advantages over RAG: first, it ensures data fidelity and completeness by granting models direct access to raw, real-time data sources; second, it shifts the model’s role from a passive question-answerer to an active system operator. To the best of our knowledge, applying a standardized protocol such as MCP to ground LLMs in real-time IoT monitoring remains a novel and underexplored research area, which our work aims to address.
@@ IV-B Inter-rater Reliability Analysis
We further performed a quantitative assessment of agreement among evaluators by computing inter-rater reliability (IRR). Inter-rater reliability measures the degree of consistency among different evaluators who assess the same set of responses [mchugh2012interrater]. In this study, IRR is crucial because it ensures that the performance scores assigned to system outputs are not biased by individual subjectivity, but instead reflect a consistent and reproducible judgment across multiple raters. The evaluation was conducted based on four major criteria: Factual Accuracy, Completeness, No Hallucination, and Usefulness. Each criterion was rated on a five-point Likert scale, where a score of 5 indicated the highest quality labeled as “Completely Accurate” and a score of 1 indicated the lowest quality labeled as “Completely Inaccurate”, with intermediate points representing varying degrees of accuracy and reliability. Six expert reviewers independently evaluated a total of 15 responses generated by the model, assigning scores for each criterion. This structured approach ensured that the assessment captured multiple dimensions of performance while providing a standardized framework for comparing judgments across raters. The aggregated results of the evaluation are presented in Table I.
@@ V Discussion
The evaluation results provide strong evidence for the system’s functional success, while also highlighting the nuanced challenges of evaluating generative AI systems. The primary finding is the co-occurrence of very high mean scores and relatively low inter-rater reliability (IRR) coefficients. While the average ratings across all four criteria were consistently high, the agreement measures such as Weighted Kappa and ICC were notably lower. This apparent contradiction is a well-known statistical artifact in reliability research, often described as the “Kappa paradox” or “ceiling effect” [feinstein1990high]. When evaluators consistently give high ratings clustered at the upper end of the scale (scores of 4 or 5), the lack of variability reduces the sensitivity of reliability coefficients, making them appear lower even in the presence of substantive consensus. Several methodological factors also contribute to this outcome. The evaluation was based on only four criteria, which, while central to system performance, do not capture the full range of qualitative variation. With such limited criteria, evaluators may apply slightly different interpretations that reduce consistency. These issues combined highlight the difficulty of applying conventional IRR metrics in settings where system outputs are of uniformly high quality.

### v3

@@ I Introduction
With the growing emphasis on healthy living, air quality monitoring has become a significant area of research. Numerous studies have shown that indoor air pollutants can significantly affect human health [18]. Particulate Matter (PM) suspended in the air, particularly PM2.5, has been proven to be directly associated with various health risks including cardiovascular diseases, respiratory infections, and even cancer [17]. Therefore, developing air quality monitoring systems that are real-time, accurate, and easy to understand is crucial for enhancing public health awareness and formulating effective environmental policies.
Existing air quality monitoring systems still face numerous challenges in practical applications. Traditional monitoring stations, while capable of providing high-precision data, have high deployment costs and limited coverage [5], resulting in insufficient spatial and temporal resolution of data [6]. More importantly, these systems primarily present data through complex charts and indicators, making it difficult for ordinary users without professional backgrounds to intuitively understand the meaning behind the data and make informed decisions accordingly. Furthermore, traditional systems generally have poor interactivity, preventing users from querying specific questions of interest in a natural and direct manner, thereby limiting the practical value of data and public engagement [19].
@@
Despite the tremendous potential of LLMs in the field of data analysis, reliably integrating them with real-time environmental monitoring data remains an under-explored research topic. Traditional LLM applications often suffer from “hallucinations” due to lack of real-time contextual knowledge, providing inaccurate or outdated information [12]. Unlike traditional retrieval-based methods, our system leverages the Model Context Protocol (MCP) to allow LLMs to securely access real-time IoT data, ensuring grounded, context-aware responses [11]. Figure 1 illustrates the overall architecture of our proposed system. Our main contributions can be summarized as follows:
@@ II Related Work
Air Quality Monitoring and Accessibility: Traditional air quality monitoring systems have matured in terms of sensor networks and data acquisition. Saini et al. provided a comprehensive overview of indoor air quality monitoring systems, documenting advances in sensor network design, pollutant detection, and data collection mechanisms [19]. These systems often achieve impressive technical performance. However, Gupta et al. shown that the information presented by many digital platforms is not always interpretable to non-expert audiences, revealing a significant gap between data acquisition and user-friendly communication of results [10]. Morawska et al. emphasized the rapid proliferation of low-cost sensors and community-driven monitoring networks, which greatly improve spatial and temporal resolution, yet, the sheer volume and heterogeneity of data can overwhelm typical interfaces, reducing their effectiveness for general users [16]. Recently, Aguado et al. evaluated low-cost indoor air quality sensors in real-world settings for CO2 and particulate matter, and while the devices supported trend tracking, the study underscores the need for intuitive feedback systems that help non-experts understand what the readings mean for health and behavior [1]. Taken together, this body of work shows that although sensor technology has matured and become more affordable, the critical challenge remains: designing interfaces that can translate dense, technical measurements into actionable insights that non-technical users can easily grasp and act upon.
@@
LLMs for Conversational Interfaces: Recent advances in LLMs have introduced new opportunities to make complex data more accessible through conversational interfaces. By allowing natural language queries, LLMs lower the barrier for non-technical users. However, In two separate studies, Dam et al. and Huang et al. both emphasized a central challenge: the propensity of LLMs for hallucination, where models produce fluent but factually incorrect responses when operating without sufficient context [7, 12]. To mitigate this, Brown et al. demonstrated that directly injecting sensor data into the model’s prompt could enable context-aware summaries and answers about specific datasets [4]. While effective at small scale, their approach is constrained by the finite context windows of current LLMs. More recently, Islam et al. proposed a prompt ensemble strategy that improved reliability of LLMs response by aggregating outputs from multiple prompts, offering a promising pathway to reduce hallucination and enhance robustness in domain-specific applications [13].
@@
Retrieval Augmented Generation for Scalability: The scalability limitations of direct context injection have been well established, most notably through the “lost-in-the-middle” phenomenon, where model performance degrades when long sequences are provided [14]. To address this issue, Gao et al. introduced Retrieval-Augmented Generation (RAG) as a dominant framework [9]. RAG operates by retrieving a targeted subset of relevant information from a larger knowledge base and passing only this filtered context to the LLM. While this approach has proven effective across a wide range of applications, Barnett et al. pointed out that its success is highly sensitive to the quality of the retrieval process, including choices around chunking strategy and parameter tuning [3]. Furthermore, RAG is not ideally suited for structured, time-series data such as environmental sensor streams, where semantic chunking fails to capture the inherent temporal dependencies in the data.
@@
Toward Agentic Systems with MCP: Building on the limitations of retrieval-based methods, our work adopts an active, agentic framework. Rather than having the system predefine which information to supply, the decision is delegated directly to the model. To enable this, Anthropic et al. introduced the Model Context Protocol (MCP), which allows LLMs to autonomously call a predefined set of secure tools to interact with backend systems [2]. This approach provides two key advantages over RAG: first, it ensures data fidelity and completeness by granting models direct access to raw, real-time data sources; second, it shifts the model’s role from a passive question-answerer to an active system operator. To the best of our knowledge, applying a standardized protocol such as MCP to ground LLMs in real-time IoT monitoring remains a novel and underexplored research area, which our work aims to address.
@@ IV-B Inter-rater Reliability Analysis
We further performed a quantitative assessment of agreement among evaluators by computing inter-rater reliability (IRR). Inter-rater reliability measures the degree of consistency among different evaluators who assess the same set of responses [15]. In this study, IRR is crucial because it ensures that the performance scores assigned to system outputs are not biased by individual subjectivity, but instead reflect a consistent and reproducible judgment across multiple raters. The evaluation was conducted based on four major criteria: Factual Accuracy, Completeness, No Hallucination, and Usefulness. Each criterion was rated on a five-point Likert scale, where a score of 5 indicated the highest quality labeled as “Completely Accurate” and a score of 1 indicated the lowest quality labeled as “Completely Inaccurate”, with intermediate points representing varying degrees of accuracy and reliability. Six expert reviewers independently evaluated a total of 15 responses generated by the model, assigning scores for each criterion. This structured approach ensured that the assessment captured multiple dimensions of performance while providing a standardized framework for comparing judgments across raters. The aggregated results of the evaluation are presented in Table I.
@@ V Discussion
The evaluation results provide strong evidence for the system’s functional success, while also highlighting the nuanced challenges of evaluating generative AI systems. The primary finding is the co-occurrence of very high mean scores and relatively low inter-rater reliability (IRR) coefficients. While the average ratings across all four criteria were consistently high, the agreement measures such as Weighted Kappa and ICC were notably lower. This apparent contradiction is a well-known statistical artifact in reliability research, often described as the “Kappa paradox” or “ceiling effect” [8]. When evaluators consistently give high ratings clustered at the upper end of the scale (scores of 4 or 5), the lack of variability reduces the sensitivity of reliability coefficients, making them appear lower even in the presence of substantive consensus. Several methodological factors also contribute to this outcome. The evaluation was based on only four criteria, which, while central to system performance, do not capture the full range of qualitative variation. With such limited criteria, evaluators may apply slightly different interpretations that reduce consistency. These issues combined highlight the difficulty of applying conventional IRR metrics in settings where system outputs are of uniformly high quality.